OpenStudy (anonymous):

2x + 3y = 1200, write in slope-intercept form?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (azureilai):

You want to isolate the y to one side 2x+3y=1200 subtract the 2x from the left to isolate y 3y=-2x+1200 I will leave the simplification up to you.

OpenStudy (anonymous):

it'd be -2/3x ? @Azureilai

OpenStudy (skullpatrol):

What does "slope-intercept form" mean?

OpenStudy (skullpatrol):

@sativamorgann Are you still there?

OpenStudy (skullpatrol):

$$\huge y=mx+b$$

OpenStudy (azureilai):

you already divided the -2 by 3, now you only need to divide 1200 by 3 to get the form. ( you divide all terms so equation remains balanced)

OpenStudy (skullpatrol):

(-2/3)x is the mx part of the form y = mx + b